Write all the factors of the following numbers: 
15??????????

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

15 = 1 × 15 15 = 3 × 5 15 = 5 × 3
Stop here since 3 and 5 have occurred earlier
Hence, the factors of 15 are 1, 3, 5 and 15

Thanks